OpenStudy (anonymous):

Could someone please help me with this problem? :c 5. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My f(x) is y=1(x-1)^2+12 Pretty Please?

OpenStudy (anonymous):

average rate of change is \[\frac{f(b)-f(a)}{b-a}\] which in your case is \[\frac{f(7)-f(4)}{7-4}\] pretty clear that \(7-4=3\) so what you need now is \(f(7)\) and \(f(4)\) so you can compute the numerator

OpenStudy (anonymous):

Functions aren't really my strong suit... Like, I have no idea what f(4) or f(7) even is .-. If not just trying to get answers though, if you could help me understand further that would be great!(: @satellite73

OpenStudy (anonymous):

I'm* @satellite73

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[f(x)=(x-1)^2+12\] right?

OpenStudy (anonymous):

that makes \[f(7)=(7-1)^2+12=6^2+12=36+12=48\]

OpenStudy (anonymous):

you good so far? we are going to do the same thing for \(f(4)\) \[f(4)=(4-1)^2+12=3^2+12=9+12=21\]

OpenStudy (anonymous):

and therefore \(f(7)-f(4)=48-21=27\)

OpenStudy (anonymous):

final answer is \(\frac{27}{3}=9\)

OpenStudy (anonymous):

Ohhhhhh! Okay wait, so does the number in the () effect the equation at all? Like I see that you use that number to substitute but, do you do anything with it other than just use it for x? @satellite73

OpenStudy (anonymous):

And how did you get the 3 under the 27? .-. @satellite73

OpenStudy (ranga):

Look at the first reply. The denominator is (b - a) = (7 - 4) = 3

OpenStudy (anonymous):

@ranga Ohhh! Thank you!!!(:

OpenStudy (ranga):

you are welcome.